OpenStudy (honeymonster):

ok..how do i make x the subject of the formula when fractions are involved :S..... equation posted in replies..

OpenStudy (honeymonster):

\[\frac{ 1 }{ x } + \frac{ 1 }{ y } = 1\]

hartnn (hartnn):

since 1/y is added,first u can subtract 1/y from both sides..what will u get?

OpenStudy (honeymonster):

1/x =y

hartnn (hartnn):

nopes,u should subtract 1/y to 1,like this \[\frac{1}{x}=1-\frac{1}{y}\] got this?

OpenStudy (honeymonster):

oh ok...so would you then multiply?

hartnn (hartnn):

multiply what?

OpenStudy (honeymonster):

the fractions to cancel them out? :S or am i going completely the wrong direction?

hartnn (hartnn):

ok,i will show u next step,see if u can understand: \[\frac{1}{x}=\frac{y-1}{y}\] got this?

OpenStudy (honeymonster):

ok i understand that bit..

hartnn (hartnn):

now since there is 1/x which is reciprocal of x, i would just invert both sides like this: \[x=\frac{y}{y-1}\] got this?

OpenStudy (honeymonster):

ok..i think i get that..so basically you would just swap the top and the bottom numbers on the right side?

hartnn (hartnn):

yes,because there was 1/x on left side and u did the same thing on both sides.

OpenStudy (honeymonster):

ok thank you :)

hartnn (hartnn):

welcome :)